### v2

@@ 3.1 FGN contracts
A contract $f$ is failures-get-nothing (FGN) if $f_{i}(S)=0$ whenever $i\notin S$.
The following result shows that, in terms of implementing any given equilibrium $p\in\mathcal{E}$, one can restrict attention to FGN contracts.
{proprep}
$\mathcal{E}=E(\mathcal{F}_{FGN})$.
@@
Let $p\in\mathcal{E}$ and let $f\in\mathcal{F}$ be a contract for which $p$ is an equilibrium. Consider the contract $g$ defined by
@@
$\lambda_{i}=\begin{cases}\dfrac{r_{i}(f,p_{-i})}{\mathbb{E}[f_{i}(S)\,|\,i\in S]}&\text{ if }p_{i}>0\\ 0&\text{ if }p_{i}=0.\end{cases}$
Notice that it follows from the first-order condition in Equation 1 that $\lambda_{i}\leq 1$ so that $g$ is indeed an FGN contract. Now if $p_{i}>0$, agent $i$’s marginal utility at profile $p$ under the contract $g$ is
$\dfrac{\partial u_{i}(p)}{\partial p_{i}}$ $=\mathbb{E}[g_{i}(S)\,|\,i\in S]-c_{i}^{\prime}(p_{i})$
$=\lambda_{i}\mathbb{E}[f_{i}(S)\,|\,i\in S]-c_{i}^{\prime}(p_{i})$
$=r_{i}(f,p_{-i})-c_{i}^{\prime}(p_{i})$
$=0,$ (because $(\text{because }p\in E(f))$
so by concavity of payoffs, $b_{i}(g,p_{-i})=p_{i}$. If $p_{i}=0$, then $g_{i}(S)=0$ for all $S$, so $b_{i}(g,p_{-i})=0=p_{i}$. Hence, $p\in E(g)$. ∎
To prove Proposition 3.1, we show that starting with any $p\in\mathcal{E}$ and any contract $f$ which implements $p$, there is a scaling factor $\lambda_{i}$ for each agent $i$ so that the FGN contract
implements $p$.
Working with FGN contracts is thus without loss of generality, and is typically easier than working with general contracts.
To begin, it allows us to work with a simpler first order condition. Observe that if $f$ is an FGN contract, then $r_{i}^{f}(p_{-i})=\mathbb{E}[f_{i}(S)\,|\,i\in S]\geq 0$, and so by Equation (1), agent $i$’s best response $b_{i}^{f}(p_{-i})$ satisfies the first order condition
$c_{i}^{\prime}(b_{i}^{f}(p_{-i}))=r_{i}^{f}(p_{-i}).$ (2)
It follows then that for any FGN contract $f$, the expected total reward at $p\in[0,1]^{n}$ to agents in $I\subset[n]$ is
$\begin{split}\mathbb{E}\left[\sum_{i\in I}f_{i}(S)\right]&=\sum_{i\in I}\mathbb{E}\left[f_{i}(S)|i\in S\right]\cdot\mathbb{P}[i\in S]\\ &=\sum_{i\in I}r_{i}^{f}(p_{-i})\cdot p_{i}\\ &=\sum_{i\in I}p_{i}\cdot c_{i}^{\prime}(b_{i}^{f}(p_{-i})).\end{split}$ (3)
In particular, if $p\in\mathcal{E}$, and $f$ is an FGN contract that implements $p$, then $b_{i}^{f}(p_{-i})=p_{i}$. Using Equation (3), we obtain the following relationship between an implementable profile $p\in\mathcal{E}$ and any contract $f\in\mathcal{F}_{FGN}$ that implements $p$.
{lemrep}
If $p\in\mathcal{E}$ and $f\in E^{-1}(p)\cap\mathcal{F}_{FGN}$, then
$\sum_{i\in[n]}{p_{i}\cdot c_{i}^{\prime}(p_{i})}=\mathbb{P}[S\neq\emptyset]\cdot\mathbb{E}\left[\sum_{i\in S}{f_{i}(S)}\,|\,S\neq\emptyset\right].$
Proof.
Since $p\in E(f)$ and $f\in\mathcal{F}_{FGN}$, it follows from Equation 3 that
$\mathbb{E}\left[\sum_{i\in[n]}f_{i}(S)\right]=\sum_{i\in[n]}p_{i}c_{i}^{\prime}(p_{i})).$
And since $f\in\mathcal{F}_{FGN}$,
$\mathbb{E}\left[\sum_{i\in[n]}f_{i}(S)\right]=\mathbb{P}[S\neq\emptyset]\cdot\mathbb{E}\left[\sum_{i\in S}{f_{i}(S)}\,|\,S\neq\emptyset\right],$
which gives the result. ∎
Moreover, since $\sum_{i\in S}{f_{i}(S)}\leq 1$, it follows from Lemma 3.1 that for any $p\in\mathcal{E}$,
$\sum_{i}{p_{i}\cdot c_{i}^{\prime}(p_{i})}\leq\mathbb{P}[S\neq\emptyset]\iff\sum_{i}{p_{i}\cdot c_{i}^{\prime}(p_{i})}+\prod_{i}(1-p_{i})\leq 1.$
Defining $z:[0,1)^{n}\to\mathbb{R}_{+}$ by
$z(p)=\sum_{i}{p_{i}\cdot c_{i}^{\prime}(p_{i})}+\prod_{i}(1-p_{i}),$
we note this necessary condition for $p$ to be implementable in the following Corollary.
Corollary 1.
If $p\in\mathcal{E}$, then $z(p)\leq 1$.
3.2 SGE contracts
As we saw in Proposition 3.1, FGN contracts are sufficient to implement any implementable equilibrium $p\in\mathcal{E}$. However, some FGN contracts always lead to non-maximal equilibria, and so would be suboptimal for any objective. For instance, a contract which never gives any reward to any agent has a unique equilibrium $p=(0,\dots,0)$, which is clearly not maximal. In this subsection, we identify a subclass of FGN contracts whose equilibria characterize the set of maximal equilibria $\mathcal{P}$.
Consider first the special case where there is only a single agent. In this case, it is easy to see that $\mathcal{P}$ contains only a single element, which is the unique equilibrium of the contract in which the agent is rewarded with the entire budget if it succeeds and nothing if it fails. Thus, in addition to not giving any reward to the agent when it fails, it is optimal to reward the agent with the entire budget if it succeeds. As our first main result shows, this property of optimal contracts generalizes to more than one agent.
@@ 3.2 SGE contracts
{thm}
@@
Suppose $p$ is an equilibrium under contract $f$. Then, $p$ is a maximal equilibrium if and only if $f$ is a successful-get-everything contract.
In other words, using SGE contracts is necessary and sufficient for implementing maximal equilibria. In particular, for any objective of the principal, any optimal contract must be SGE.
@@
Observe that contracts like a piece-rate contract (where each agent $i$ gets a fixed reward if she is successful and 00 otherwise), or a bonus pool contract (where each agent gets a nonzero share of the budget only if all agents succeed), are not successful-get-everything. It follows from Theorem 3.2 that such contracts never implement maximal equilibria, and so the choice of any such contract would be strictly suboptimal for any objective. In comparison, SGE contracts induce competition between the agents as they compete for a fixed budget, and from Theorem 3.2, always lead to maximal equilibria. Thus, the result suggests that in environments where the principal operates with an exclusive-use budget, fostering competition among the agents through successful-get-everything contracts creates stronger incentives than promoting teamwork through bonus-pool contracts or independent performance assessment through piece-rate contracts.
3.3 Application: two agents, quadratic costs
By illustrating the suboptimality of contracts that award agents when they fail or do not always exhaust the entire budget, Theorem 3.2 also reduces the complexity of the contract design problem. We illustrate this by solving the contract design problem for the special case where there are two agents with quadratic costs, $c_{i}(p_{i})=\frac{1}{2}C_{i}p_{i}^{2}$.
With two agents, while a typical contract has eight parameters, since each agent’s reward must be specified for each possible outcome, the set of SGE contracts is parameterized by a single parameter $\lambda$ that represents agent 1’s share of the budget when both agents succeed, $\lambda=f_{1}(\{1,2\})$. To see why, note that $f_{i}(S)=0$ if $i\notin S$ and $f_{i}(S)=1$ if $S=\{i\}$, so the only free variables are $f_{i}(\{1,2\})$, and since $f_{2}(\{1,2\})=1-f_{1}(\{1,2\})$, $f$ is entirely determined by $f_{1}(\{1,2\})$. The utilities of the agents under $f$ are
$u_{1}(p)$ $=p_{1}\cdot(1-p_{2})\cdot 1+p_{1}\cdot p_{2}\cdot\lambda-\frac{1}{2}C_{1}p_{1}^{2}$
$u_{2}(p)$ $=(1-p_{1})\cdot p_{2}\cdot 1+p_{1}\cdot p_{2}\cdot(1-\lambda)-\frac{1}{2}C_{2}p_{2}^{2}$
and using Equation (2), it follows that the unique equilibrium under $f$ is
$p_{1}(\lambda)=\frac{C_{2}-(1-\lambda)}{C_{1}C_{2}-\lambda(1-\lambda)}\quad\quad\quad p_{2}(\lambda)=\dfrac{C_{1}-\lambda}{C_{1}C_{2}-\lambda(1-\lambda)}\cdot$
By Theorem 3.2, this provides the set of maximal equilibria,
$\mathcal{P}=\left\{\big{(}p_{1}(\lambda),p_{2}(\lambda)\big{)}\,:\,\lambda\in[0,1]\right\}.$
The principal’s problem is thus to find the $\lambda$ that maximizes $V(p_{1}(\lambda),p_{2}(\lambda))$. In the following result, we consider the particular case where $V$ is linear, and describe how $\lambda$ might depend on the heterogeneity between the agents, and the principal’s own bias. {thmrep} Suppose $n=2$, $c_{i}(p_{i})=\frac{1}{2}C_{i}p_{i}^{2}$ with $C_{i}>1$, and $V(p_{1},p_{2})=wp_{1}+\cdot p_{2}$. Then, the optimal contract, defined by $\lambda_{1}(w)$, takes the form
@@ 3.3 Application: two agents, quadratic costs
$\lambda_{1}(w)=\begin{cases}0,&\text{ if }w\leq\frac{C_{1}C_{2}-C_{1}}{C_{1}C_{2}+C_{2}-1}\\ \frac{1}{2},&\text{ if }w=1\\ 1,&\text{ if }w\geq\frac{C_{1}C_{2}+C_{1}-1}{C_{1}C_{2}-C_{2}}\end{cases}.$

### v3

@@ 3 Main results
A contract $f\in\mathcal{F}$ is failures-get-nothing (FGN) if $f_{i}(S)=0$ whenever $i\notin S$.
It turns out that using a contract which is not FGN is wasteful in a very particular sense. Given any such contract and an effort profile $p$ it implements, it is possible to scale the payments down at every outcome to form an FGN contract which also implements $p$. As a consequence, FGN contracts are not only sufficient, but offer the cheapest alternative for inducing any implementable profile. {proprep} For any implementable profile $p$, there exists an FGN contract $g$ which implements it. Moreover, the expected total payment at $p$ under $g$ is minimal:
$\mathbb{E}_{p}[\sum_{i\in[n]}g_{i}(S)]\cdot B=\sum_{i\in[n]}p_{i}c_{i}^{\prime}(p_{i}).$
@@
Let $p\in\mathcal{E}$ and let $f\in E^{-1}(p)$. Consider the contract $g$ defined by
@@
$\lambda_{i}=\begin{cases}\dfrac{\mathbb{E}_{p}[f_{i}(S)\,|\,i\in S]-\mathbb{E}_{p}[f_{i}(S)\,|\,i\notin S]}{\mathbb{E}_{p}[f_{i}(S)\,|\,i\in S]}&\text{ if }p_{i}>0\\ 0&\text{ if }p_{i}=0.\end{cases}$
It follows that $\lambda_{i}\leq 1$ so that $g$ is indeed an FGN contract. Now if $p_{i}>0$, agent $i$’s marginal utility at profile $p$ under the contract $g$ is
[…]
$=\lambda_{i}\mathbb{E}_{p}[f_{i}(S)\,|\,i\in S]\cdot B-c_{i}^{\prime}(p_{i})$
$=\mathbb{E}_{p}[f_{i}(S)\,|\,i\in S]\cdot B-\mathbb{E}_{p}[f_{i}(S)\,|\,i\notin S]\cdot B-c_{i}^{\prime}(p_{i})$
$=0.$ (because $(\text{because }p\in E(f))$
By concavity of payoffs, $b_{i}(g,p_{-i})=p_{i}$. If $p_{i}=0$, then $g_{i}(S)=0$ for all $S$, so $b_{i}(g,p_{-i})=0=p_{i}$. Hence, $p\in E(g)$. It is straightforward to verify that the expected total payment under $g$ at $p$ is $\sum_{i=1}^{n}p_{i}c_{i}^{\prime}(p_{i})$, and it is weakly lower than any other $f$ that may induce $p$. ∎
Going forward, without loss of generality, we restrict attention to FGN contracts. Under any FGN contract $f$ and effort profile $p$, Equation (2) simplifies to:
$c_{i}^{\prime}(p_{i})=\mathbb{E}_{p}[f_{i}(S)\mid i\in S]\cdot B.$ (3)
This equation shows clearly that an agent’s incentive to exert effort increases in their expected share of the budget if they succeed. Since the principal’s budget is exclusive-use, meaning that it does not derive any value from leftover budget, this suggests that the principal should always distribute the entire budget among successful agents.
@@
Given any implementable profile $p$, if the FGN contract $f$ which induces it is not SGE, there must be some leftover budget under some outcomes. Observe that the most the principal can give in expected rewards at $p$ is $\mathbb{P}_{p}[S\neq\emptyset]\cdot B$, and the actual amount it gives in expected rewards is $\mathbb{E}_{p}[\sum_{i}f_{i}(S)]\cdot B$. We denote by $z(p)$ this gap between the upper bound and the actual amount given in rewards, and note that it depends only on $p$:
$z(p)$ $=\mathbb{P}_{p}[S\neq\emptyset]\cdot B-\mathbb{E}_{p}\left[\sum_{i}f_{i}(S)\right]\cdot B$
$=\left[1-\left(\prod_{i\in[n]}(1-p_{i})\right)\right]\cdot B-\sum_{i\in[n]}p_{i}c_{i}^{\prime}(p_{i}),$
where the second term follows from Proposition 2. For any $p\in\mathcal{E}$, $z(p)$ provides a useful robust measure of the leftover budget from inducing $p$ using any FGN contract $f\in E^{-1}(p)$.
Corollary 1.
If $p$ is implementable, then $z(p)\geq 0$. Moreover, if $p$ is implementable by an SGE contract, then $z(p)=0$.
Now suppose effort profile $p$ is an equilibrium of a SGE contract. If $q$ is an effort profile that coordinate-wise dominates $p$, we can show that $z(q)<z(p)=0$, and so by Corollary 1, $q$ is not implementable. It thus follows that any equilibrium of any SGE contract is maximal.
Interestingly, maximal equilibria can only be induced by SGE contracts. Intuitively, if $p$ is an an equilibrium of a FGN contract which is not SGE, the leftover budget could potentially be used to incentivize higher effort. While one might suspect that increasing agent $i$’s share $f_{i}(S)$ for some $i\in S$ (if $\sum_{i\in S}f_{i}(S)<1$) might do the trick, the equilibrium forces imply that such a transformation does not necessarily lead to a dominant outcome. [^5] For an illustration, consider an example with $n=2$, $c_{i}(p_{i})=\frac{1}{2}C_{i}p_{i}^{2}$, and $B=1$. Suppose $f\in\mathcal{F}$ is such that it awards $\beta_{i}\in[0,1]$ to agent $i$ if and only if $i$ is the only successful agent. Using Equation (3), one can verify that $f$ has a unique equilibrium $p=(p_{1},p_{2})$ with $p_{i}=\dfrac{\beta_{i}(C_{-i}-\beta_{-i})}{C_{1}C_{2}-\beta_{1}\beta_{2}}.$ From here, it is easy to check that increasing $\beta_{i}$ would increase $p_{i}$, but it may also be accompanied by a decrease in $p_{-i}$. In general, even though there isn’t an obvious way to transform these contracts so that they induce dominant outcomes, we show there always exists some way.
{thmrep}
@@
Suppose the contract $f$ implements the effort profile $p$. Then, $p$ is maximal if and only if $f$ is an SGE contract.
Proof.
Suppose $f$ is a SGE contract and $p\in E(f)$. We want to show that $p$ is maximal. Observe that
$\dfrac{\partial z}{\partial x_{i}}\Bigr{|}_{x\geq p}$ $=B\prod_{j\neq i}(1-x_{j})-x_{i}c_{i}^{\prime\prime}(x_{i})-c_{i}^{\prime}(x_{i})$
$<B\prod_{j\neq i}(1-x_{j})-c_{i}^{\prime}(x_{i})$ (because $(\text{because }c_{i}\text{ is convex})$
$\leq B\prod_{j\neq i}(1-x_{j})-c_{i}^{\prime}(p_{i})$ (because $(\text{because }c_{i}\text{ is convex})$
$\leq B\prod_{j\neq i}(1-p_{j})-c_{i}^{\prime}(p_{i})$ (because $(\text{because }x_{j}\geq p_{j})$
$\leq B\cdot\mathbb{E}_{p}[f_{i}(S)|i\in S]-c_{i}^{\prime}(p_{i})$ (because $(\text{because }f\in\mathcal{F}_{SGE})$
$=0$ (because $(\text{because }p\in E(f))$
It follows that if effort profile $q$ dominates $p$, then $z(q)<z(p)$, and from Corollary 1, $q$ is not implementable. Thus, $p$ is maximal.
Now suppose $p$ is maximal. We want to show that any $f\in E^{-1}(p)$ must be SGE. We first show that all agents must be active at $p$.
{lem}
For all $i\in[n]$, $10<p_{i}<$ 1. {nestedproof} Suppose towards a contradiction that $p_{k}=0$ for some $k\in[n]$. Let $f\in E^{-1}(p)\cap\mathcal{F}_{FGN}$ and consider the contract $g$ where $g_{i}(S)=f_{i}(S\setminus\{k\})$ for $i\neq k$ and
It is easy to verify that $p^{\prime}=(p_{k}^{\prime},p_{-k})$ where $p_{k}^{\prime}=b_{k}(g,p_{-k})>0$ is an equilibrium under $g$. But then, $p^{\prime}\in\mathcal{E}$ dominates $p$, which contradicts the fact that $p$ is maximal.
Now let
and let
$\kappa_{p}:=\{i\in[n]\,:\,\{i\}\in K_{p}\}.$
We now show that $K_{p}=\{\emptyset\}$.
  1. 1.
    Step 1: Suppose $S\in K_{p}$. For any $T\subset S$, $T\in K_{p}$.
    Let $f\in E^{-1}(p)\cap\mathcal{F}_{FGN}$ be such that $\sum_{i\in S}{f_{i}(S)}<1$. If $\sum_{i\in T}{f_{i}(T)}<1$, we are done. Otherwise, pick an agent $i\in T$ such that $f_{i}(T)>0$ and consider a contract $g$ which differs from $f$ only in its award for agent $i$ at $S$ and $T$. In particular, let $g$ be such that $g_{i}(S)=f_{i}(S)+\epsilon$ and $g_{i}(T)=f_{i}(T)-\delta$ where $\epsilon,\delta>0$ are chosen so that $p\in E(g)$. We can do this because we know from Lemma 3 that all events occur with positive probability at profile $p$. It follows then that $g\in E^{-1}(p)\cap\mathcal{F}_{FGN}$ and $\sum_{i\in T}{g_{i}(T)}<1$. Thus, $T\in K_{p}$.
  2. 2.
    Step 2: Suppose $S,T\in K_{p}$. Then, $S\cup T\in K_{p}$.
    Let $f,g\in E^{-1}(p)\cap\mathcal{F}_{FGN}$ be such that $\sum_{i\in S}{f_{i}(S)}<1$ and $\sum_{i\in T}{g_{i}(T)}<1$. Consider the contract $h=\frac{1}{2}f+\frac{1}{2}g$. From Equation (3), $h\in E^{-1}(p)\cap\mathcal{F}_{FGN}$ and also $\sum_{i\in S}{h_{i}(S)}<1$ and $\sum_{i\in T}{h_{i}(T)}<1$. Now, if $\sum_{i\in S\cup T}{h_{i}(S\cup T)}<1$, we are done. Otherwise, pick any agent $i\in S\cup T$ (WLOG, let $i\in S$) such that $h_{i}(S\cup T)>0$ and consider a contract $h^{\prime}$ which differs from $h$ only in its award for agent $i$ at $S\cup T$ and $S$. In particular, let $h^{\prime}$ be such that $h^{\prime}_{i}(S\cup T)=h_{i}(S\cup T)-\epsilon$ and $h^{\prime}_{i}(S)=h_{i}(S)+\delta$ where $\epsilon,\delta>0$ are chosen so that $p\in E(h^{\prime})$. It follows then that $h^{\prime}\in E^{-1}(p)\cap\mathcal{F}_{FGN}$ and $\sum_{i\in S\cup T}{h^{\prime}_{i}(S\cup T)}<1$. Thus, $S\cup T\in K_{p}$.
    Note that it follows from Steps 1 and 2 that $K_{p}=2^{\kappa_{p}}$.
  3. 3.
    Suppose towards a contradiction that there is an $S\subset[n]$ such that $\kappa_{p}^{C}\cap S\neq\phi$ and $f_{i}(S)>0$ for some $i\in\kappa_{p}$. Let $g\in E^{-1}(p)\cap\mathcal{F}_{FGN}$ be such that $g_{i}(\{i\})<1$. Consider the contract $h=\frac{1}{2}f+\frac{1}{2}g$. Then $h\in E^{-1}(p)\cap\mathcal{F}_{FGN}$ and also $h_{i}(S)>0$ and $h_{i}(\{i\})<1$. Now, consider a contract $h^{\prime}$ which differs from $h$ only in its award for agent $i$ at $S$ and $\{i\}$. In particular, let $h^{\prime}$ be such that $h^{\prime}_{i}(\{i\})=h_{i}(\{i\})+\epsilon$ and $h^{\prime}_{i}(S)=h_{i}(S)-\delta$ where $\epsilon,\delta>0$ are chosen so that $p\in E(h^{\prime})$. It follows then that $h^{\prime}\in E^{-1}(p)\cap\mathcal{F}_{FGN}$ and $\sum_{i\in S}{h^{\prime}_{i}(S)}<1$. But this means that $S\subset\kappa_{p}$ which is a contradiction.
  4. 4.
    Step 4: Suppose $\kappa_{p}\neq\phi$. Then there is a $p^{\prime}\in\mathcal{E}$ that dominates $p$.
    by averaging over the outcomes of agents in $\kappa_{p}$ under $g$.
    $c_{i}^{\prime}(p^{\prime}_{i})=B\cdot t_{i}(\epsilon)\cdot\sum_{S\subset A_{-i}}(h_{i}(S\cup\{i\})){\Pr}^{A_{-i}}_{p^{\prime}_{-i}}(S)$
    Observe that as $\epsilon\to 0$, $t_{i}(\epsilon)\to 1$ for all $i\in A$. Since $\sum_{i\in S}{h_{i}(S)}<1$ for all $S\subset A$ and $t_{i}(\epsilon)$ is continuous in $\epsilon$, we can find $\epsilon>0$ small enough so that the contract $h^{\prime}_{i}(S)=h_{i}(S)*t_{i}(\epsilon)$ for all $S\subset A$ and $i\in S$ is a feasible contract. By construction, the profile $(p^{\prime},p_{-A})$ will be an equilibrium under $h^{\prime}$. Thus, we have that $p$ is not maximal.
It follows then that $K_{p}=\{\emptyset\}$. By definition of $K_{p}$, this means that for any $f\in E^{-1}(p)$, either $f\notin\mathcal{F}_{FGN}$ or $f\in\mathcal{F}_{SGE}$. Now suppose there exists an $f$ such that $f\notin\mathcal{F}_{FGN}$ and $p\in E(f)$. From Proposition 2, we can find $g\in\mathcal{F}_{FGN}$ such that $p\in E(g)$. Moreover, we know from the construction of $g$ in the argument of Proposition 2 that $g\notin\mathcal{F}_{SGE}$. Thus, we have that $g\in\mathcal{F}_{FGN},g\notin\mathcal{F}_{SGE},$ and $p\in E(g)$. But this means that $K_{p}\neq\emptyset$ which is a contradiction. Therefore, it must be that for any $f\in E^{-1}(p)$, $f\in\mathcal{F}_{SGE}$.
@@
As a consequence, since contracts like piece-rate (where each agent $i$ gets a fixed reward if she is successful and 00 otherwise) or bonus-pool (where each agent gets a nonzero share of the budget only if all agents succeed) are not SGE, they never induce maximal equilibria. In comparison, SGE contracts induce competition between the agents as they compete for a fixed budget, and always lead to maximal equilibria. Thus, the result suggests that in environments where the principal operates with an exclusive-use budget, fostering competition among the agents through successful-get-everything contracts creates stronger incentives than promoting teamwork through bonus-pool contracts or independent performance assessment through piece-rate contracts.
While these simpler, well-known contracts should be avoided in favor of SGE contracts, the set of all SGE contracts itself lacks clear structure, being $\Theta(n2^{n})$-dimensional. Consequently, optimizing over SGE contracts may still pose a computationally difficult problem, and the resulting optimal contract may be complex to interpret and implement in practice. Now, the set of maximal equilibria lies on the boundary of the equilibrium set and is thus at most $(n-1)$-dimensional. This dimensionality gap suggests that for any given maximal equilibrium, there are potentially many SGE contracts that implement it, which raises the question: is there a simpler, low-dimensional, structured subclass of SGE contracts that the principal can use? Weighted contracts, which are both structured and dimensionally aligned, provide a natural and promising candidate.
Definition 4.
A weighted contract assigns weights to the agents, and distributes the entire budget among successful agents in proportion to their weights. As the following result shows, these simple contracts do in fact essentially span the whole set of maximal equilibria.
{thmrep}
For every maximal equilibrium $p$, there is a unique contract in the closure of weighted contracts which implements $p$.
Proof.
From Theorem 3, we know that $p$ must be an equilibrium of a SGE contract. But then, it follows that $p$ must satisfy (5). From Proposition 4, there is a Luce contract that implements $p$. Thus, we get that every maximal equilibrium can be implemented by a Luce contract.
Now we show that there is a unique Luce contract that implements $p$. Consider the case where $f,g$ are weighted contracts and suppose towards a contradiction that $p\in E(f)\cap E(g)$. Let $i\in[n]$ denote the agent with the smallest weight ratio $\frac{\lambda^{g}_{i}}{\lambda_{i}^{f}}$, where $\lambda^{f},\lambda^{g}$ represent the weights that define the contracts $f,g\in\mathcal{F}_{W}$. It follows from the definition of weighted contracts that for any $S\subset[n]$ such that $i\in S$, agent $i$’s reward is weakly lower under $g$ than $f$. As a result, it can’t be that $b_{i}(g,p_{-i})=b_{i}(f,p_{-i})$. It follows that two different weighted contracts cannot have the same equilibrium.
Moreover, for any $j\in T^{\prime}\setminus T$,
Thus,
However, since $p$ is an equilibrium for $f$ and $g$, both of the outside expressions must be equal to $\frac{\sum_{k\in T}p_{k}\cdot c_{k}^{\prime}(p_{k})}{B}$, contradiction. ∎
The closure of weighted contracts also has a relatively simple structure. We refer to these contracts, containing weighted contracts and their limit points, as Luce contracts.
Definition 5.
[…]
Luce contracts reward each highest-priority successful agent with a fraction of the budget proportional to her weight. [^6] The choice of terminology is motivated by the analogy between the structure of Luce contracts and the selection rules that satisfy Luce’s choice axiom (Luce (1959)). Just as Luce rules have the property that the relative odds of selecting one item over another from a pool is unaffected by which other items are in the pool, Luce contracts have the property that the reward of a successful agent relative to another successful agent is independent of which other agents succeed.
{lemrep}
A contract $f$ is in the closure of weighted contracts if and only if it is a Luce contract.
Proof.
To begin, suppose that $f$ is the limit of a sequence $g^{m}$ of weighted contracts with weights $(\lambda_{1}^{m},\dots,\lambda_{n}^{m})$. Define the relation $\succeq$ by $i\succeq j$ if and only if $f_{i}(\{i,j\})>0$. It is immediate that $\succeq$ is complete. Moreover, since $f_{i}(\{i,j\})>0$ if and only if there exists an $\varepsilon>0$ such that $\lambda_{i}^{m}\geq\varepsilon\lambda_{j}^{m}$ for all $m$, $\succeq$ is also transitive, and hence it is a non-strict order. For each $i$, let $S^{i}=\{j\,:\,i\sim j\}$, and define $\lambda_{i}=f_{i}(S^{i})$. We claim that $f$ is a Luce contract with priority ordering given by $\succeq$ and weights given by $(\lambda_{i})$. To prove this, fix $i$ and $S$.
If $i\notin S$, then
$f_{i}(S)=\lim_{m\to\infty}g^{m}_{i}(S)=\lim_{m\to\infty}0=0.$
If $i\in S\setminus Top_{\succeq}(S)$, then there is some $j\in S$ such that $j\succ i$. Now,
$g^{m}_{i}(S)=\frac{\lambda_{i}^{m}}{\sum_{k\in S}\lambda_{k}^{m}}\leq\frac{\lambda_{i}^{m}}{\lambda_{i}^{m}+\lambda_{j}^{m}}=g^{m}_{i}(\{i,j\}),$
so
$f_{i}(S)=\lim_{m\to\infty}g^{m}_{i}(S)\leq\lim_{m\to\infty}g^{m}_{i}(\{i,j\})=f_{i}(\{i,j\})=0.$
Finally, observe that for any $i$ and $j$, if $i\succ j$, then
$\lim_{m\to\infty}\frac{1}{1+\frac{\lambda_{j}^{m}}{\lambda_{i}^{m}}}=\lim_{m\to\infty}\frac{\lambda_{i}^{m}}{\lambda^{i}_{m}+\lambda_{j}^{m}}=\lim_{m\to\infty}g^{m}_{i}(\{i,j\})=f_{i}(\{i,j\})=1,$
so $\lim_{m\to\infty}\frac{\lambda_{j}^{m}}{\lambda_{i}^{m}}=0$, and if $i\sim j$, then $S^{i}=S^{j}$, so
$\lim_{m\to\infty}\frac{\lambda_{j}^{m}}{\lambda_{i}^{m}}=\lim_{m\to\infty}\frac{\frac{\lambda_{j}^{m}}{\sum_{k\in S^{j}}\lambda_{k}^{m}}}{\frac{\lambda_{i}^{m}}{\sum_{k\in S^{i}}\lambda_{k}^{m}}}=\lim_{m\to\infty}\frac{g^{m}_{j}(S^{j})}{g^{m}_{i}(S^{i})}=\frac{f_{j}(S^{j})}{f_{i}(S^{i})}=\frac{\lambda_{j}}{\lambda_{i}}\cdot$
Hence, if $i\in Top_{\succeq}(S)$, then
$f_{i}(S)$ $=\lim_{m\to\infty}g^{m}_{i}(S)=\lim_{m\to\infty}\frac{\lambda^{m}_{i}}{\sum_{k\in S}\lambda^{m}_{k}}=\lim_{m\to\infty}\frac{1}{\sum_{k\in Top_{\succeq}(S)}\frac{\lambda^{m}_{k}}{\lambda^{m}_{i}}+\sum_{k\in S\setminus Top_{\succeq}(S)}\frac{\lambda^{m}_{k}}{\lambda^{m}_{i}}}$
$=\frac{1}{\sum_{k\in Top_{\succeq}(S)}\frac{\lambda_{k}}{\lambda_{i}}}=\frac{\lambda_{i}}{\sum_{k\in Top_{\succeq}(S)}\lambda_{k}}\cdot$
Since $f$ was arbitrary, it follows that $\bar{F}_{W}\subseteq\mathcal{F}_{Luce}$.
Now, suppose $f$ is a Luce contract. For each $i$, let $r_{i}=|\{j\,:\,j\succeq i\}|$, and for every $t>0$, let $g^{t}$ be the weighted contract with weights given by $\lambda^{t}_{i}=t^{r_{i}}\cdot\lambda_{i}$. Now, fix $i$ and $S$. If $i\notin S$,
$\lim_{t\to 0}g^{t}_{i}(S)=\lim_{t\to 0}0=0=f_{i}(S).$
If $i\in S\setminus Top_{\succeq}(S)$, then there is some $j\in S$ such that $j\succ i$ and hence $r_{j}<r_{i}$. In particular,
$\lim_{t\to 0}g^{t}_{i}(S)=\lim_{t\to 0}\frac{t^{r_{i}}\cdot\lambda_{i}}{\sum_{k\in S}t^{r_{k}}\cdot\lambda_{k}}\leq\lim_{t\to 0}\frac{t^{r_{i}}\lambda_{i}}{t^{r_{j}}\lambda_{j}}=0=f_{i}(S).$
Finally, if $i\in Top_{\succeq}(S)$, then
$g^{t}_{i}(S)=\frac{t^{r_{i}}\lambda_{i}}{\sum_{k\in S}t^{r_{k}}\cdot\lambda_{k}}$ $=\frac{t^{r_{i}}\lambda_{i}}{t^{r_{i}}\cdot\sum_{k\in Top_{\succeq}(S)}\lambda_{k}+\sum_{k\in S\setminus Top_{\succeq}(S)}t^{r_{k}}\cdot\lambda_{k}}$
$=\frac{\lambda_{i}}{\sum_{k\in Top_{\succeq}(S)}\lambda_{k}+\sum_{k\in S\setminus Top_{\succeq}(S)}t^{r_{k}-r_{i}}\cdot\lambda_{k}}.$
Since $r_{k}>r_{i}$ for all $k\in S\setminus Top_{\succeq}(S)$, it follows that
$\lim_{t\to 0}g^{t}_{i}(S)=\frac{\lambda_{i}}{\sum_{k\in Top_{\succeq}(S)}\lambda_{k}}=f_{i}(S).$
Thus, $f=\lim_{t\to 0}g^{t}$. Since $f$ was arbitrary, it follows that $\mathcal{F}_{Luce}\subseteq\bar{\mathcal{F}}_{W}$. ∎
It follows that the principal can always just optimize its objective over the class of Luce contracts.
Corollary 2.
For every maximal equilibrium $p$, there is a unique Luce contract which implements $p$. Thus, for any strictly increasing continuous objective $V(p)$,
Despite this reduction in complexity, it remains difficult to explicitly solve for the optimal Luce contract for a given objective in general. We are able to solve the problem for the special case where there are two agents with quadratic costs.
{proprep}
Suppose $n=2$, $c_{i}(p_{i})=\frac{1}{2}C_{i}p_{i}^{2}$ with $C_{i}>B$, and $V(p_{1},p_{2})=wp_{1}+\cdot p_{2}$. Then, the optimal contract, defined by $\lambda_{1}(w)$, takes the form
@@
$\lambda_{1}(w)=\begin{cases}0,&\text{ if }w\leq\frac{C_{1}C_{2}-BC_{1}}{C_{1}C_{2}+BC_{2}-B^{2}}\\ \frac{1}{2},&\text{ if }w=1\\ 1,&\text{ if }w\geq\frac{C_{1}C_{2}+BC_{1}-B^{2}}{C_{1}C_{2}-BC_{2}}\end{cases}.$